# Balancing Intensity and Focality in Directional DBS Under Uncertainty: A Simulation Study of Electrode Optimization via a Metaheuristic L1L1 Approach

### v1

@@ 4 Discussion
This study explores the application of the recently introduced metaheuristic L1L1 method [32], to optimize the current patterns for DBS leads. The core focus lies in L1L1’s capacity to incorporate a priori uncertainty through an explicit constraint, offering a notable advantage over traditional approaches, which often fail to account for such variability [37, 38, 36, 59]. Specifically, we define the uncertainty through a threshold on the nuisance current density, which is determined by the relative attenuation of the lead field with respect to a target site. This attenuation is quantified in decibels and corresponds to typical falloff patterns observed in VTA. By reducing the dependence on conventional parameter tuning methods, such as trial-and-error [60, 61], this approach paves the way for more sophisticated techniques to make more informed decisions, even in the presence of inherent uncertainties in DBS settings and simulations [62, 63, 64, 65].
Optimization was carried out using three parameters: the focused current density $\Gamma$, the nuisance current density $\Xi$, and field ratio $\Theta=\Gamma/\Xi$. To assess robustness and focality under uncertainty, the recently introduced L1L1 method was applied in two variants, L1L1(A) and L1L1(B), differing in the range of the nuisance current threshold $\varepsilon$, which reflects varying levels of a priori modeling uncertainty. Specifically, L1L1(B) constrains $\varepsilon$ to the interval $[-10,0]$ dB, accounting for higher uncertainty, whereas L1L1(A) utilizes a broader range $[-160,0]$ dB. For comparison, the Reciprocity Principle (RP) and Tikhonov-Regularized Least Squares (TLS) methods were included as reference approaches.
The results reveal that the spatial distributions of $\Gamma$, $\Xi$, and $\Theta$ are strongly influenced by both the optimization method and the orientation of the dipole current. Parallel configurations resulted in highly localized activation near the tip of the lead, whereas perpendicular configurations yielded a broader, medially shifted targeting. A comparative analysis across varying noise levels demonstrated that L1L1 methods consistently outperformed RP and TLS in maintaining higher and more stable values of $\Theta$, especially at higher PSNR values.
In noiseless conditions, solutions obtained with a low nuisance threshold $\varepsilon\in[-160,0]\,\text{dB}$, as in L1L1(A), often exhibited unrealistically high field ratio values $\Theta$, sometimes up to fifty times greater than those derived using the more conservative threshold $\varepsilon\in[-10,0]\,\text{dB}$ in L1L1(B). This tendency suggests overfitting when the model is not constrained by uncertainty. Introducing this noise regularized the solution space, preventing the algorithm from converging to over-idealized solutions that mirrored the original model structure. These findings underscore the significance of explicitly modeling uncertainty and justify the constraint imposed on the nuisance current in L1L1(B), which leads to more realistic and generalizable solutions. Overall, these observations highlight the potential of L1L1(B) as a robust and uncertainty-aware optimization strategy for dDBS lead configuration.
@@
The estimates for the set $\mathcal{V}_{\delta}$, where the lead field amplitude attenuates by a factor $\delta$ for $\delta\in[-40,-10]$ dB, align closely with established knowledge regarding VTA. Specifically, for $\delta\approx-40\,\text{dB}$, our model predicts attenuation within a spatial region consistent with a VTA radius of approximately 4.0 mm, in line with the findings of Butson’s [33] and McIntyre’s [20]. This agreement suggests that the noise levels used, with PSNR values of 40 dB and 50 dB, represent reasonable bounds for physiological uncertainty. At 40 dB, the dynamic range of $\Theta$ is determined by the maximum lead field attenuation within the VTA, while 50 dB further supports the assumption that each position within the VTA can be independently and coherently targeted. In this context, the L1L1 method, particularly L1L1(B), demonstrated superior targeting capabilities compared to RP and TLS, especially under noisy conditions.
@@
When comparing the electrode configurations, increasing the number of contact points generally enhances both adaptability and control over the resulting electric field. This improvement is driven by several factors, including reduced distances between electrodes and target locations within the lead field matrix, better alignment of the induced electric field with the target current direction, and a greater capacity for flexible fitting due to the increased number of degrees of freedom.
@@
The findings of this study provide a proof-of-concept for integrating a priori uncertainty modeling into numerically simulated dDBS optimization. When relevant sources of uncertainty, such as anatomical variability, conductivity inhomogeneities, or lead placement deviations, are known or can be estimated, the L1L1 method might be employed to improve the focality and directionality of stimulation, enabling more selective targeting of desired structures while reducing activation in surrounding regions. In clinical practice, such improvements might lead to more personalized and effective treatment strategies, enhancing therapeutic precision, and minimizing side effects [10].
@@ 4.1 Limitations and Future Work
While the current uncertainty modeling strategy provides a principled framework for constraining the optimization based on lead field attenuation, it also has several limitations. First, the threshold values used to define acceptable nuisance current levels are heuristic and may not fully capture the complexity of individual patient anatomies or tissue conductivities. This simplification assumes spatial uniformity in attenuation, which may not hold, for example, in the presence of anisotropic or heterogeneous tissue properties [44, 66]. Additionally, the approach relies on a static surrogate model of the VTA and does not directly incorporate a probabilistic model of uncertainty [62], which would allow for a more comprehensive treatment of variability. Potential alternative surrogate approaches exist, such as, threshold-based curve fitting [67], artificial neural networks [68], and statistical emulators such as Gaussian process classifiers [69, 70], which approach VTA in different ways.
Compared to the reference methods of this study, the L1L1 method is also slightly more demanding in terms of computing time. However, the increase is not significant and does not limit its practical use.
Other mathematical optimization techniques, such as Tikhonov (TLS) [38, 37], may offer promising alternatives alongside the L1L1 method for further refining the a priori uncertainty modeling adopted in this study. A logical direction for future research is the development of a posteriori models, where the current approach could be validated in more advanced experimental or clinical settings. This would enable a deeper understanding of the potential hyperparameter realizations in real-world applications. Translating these computational findings into clinical practice will require rigorous validation through empirical testing and clinical trials.
An ongoing work is the inclusion of more patient-specific anatomical and physiological data, e.g., a thalamic segmentation, to ensure more personalized and accurate stimulation strategies. Therefore, we explore the use of FE-based discretizations incorporating thalamic subdivisions via boundary-fitted multicompartment FE meshing approach and a complete electrode model, as utilized in this study.
@@ 5 Conclusions
This study represents an advancement in optimizing the volumetric current fields for DBS by leveraging the metaheuristic L1L1 method. Previous studies on contact selection have used convex optimization and patient-specific models to determine optimal stimulation settings. These methods perform well when all input parameters are fixed. However, they do not account for uncertainties in factors such as patient anatomy, lead placement, and tissue properties, which often vary in clinical practice. In principle, our L1L1-based method can address these uncertainties. Its main advantage is the ability to target the desired area while avoiding overfitting to noisy data.
Our findings demonstrate the method’s ability to refine electrode contact point configurations under varying levels of a priori lead field uncertainty. By achieving relatively high field ratios with a carefully chosen $\varepsilon$ value, the L1L1 method shows promise in enhancing the precision and effectiveness of DBS stimulation within uncertainty constraints. This, in turn, has the potential to improve individualized stimulation outcomes. An ongoing work is to integrate patient-specific data, such as thalamic segmentation, to improve stimulation accuracy. This involves using boundary-fitted multicompartment FE meshes and a complete electrode model, as in this study.
@@ 7 Author Contributions
FGP: Conceptualization, Formal analysis, Investigation, Methodology, Software, Visualization, Writing—original draft. AL: Conceptualization, Data curation, Investigation, Methodology, Project administration, Writing—original draft. MS: Formal analysis, Supervision, Validation, Writing—review & editing. SP: Conceptualization, Formal analysis, Funding acquisition, Investigation, Methodology, Project administration, Resources, Software, Supervision, Validation, Visualization, Writing—review & editing.
@@ 8 Funding
FGP, MS, and SP were supported by the Research Council of Finland through the Center of Excellence in Inverse Modelling and Imaging 2018–2025 (353089); the researcher exchange (DAAD) project ”Non-invasively reconstructing and inhibiting activity in focal epilepsy” (354976, 367453); the ERA PerEpi project (344712), and Flagship of Advanced Mathematics for Sensing, Imaging and Modelling (FAME) (359185, 359198)
@@ Appendix A Mathematical Optimization Scheme
We consider a weighted optimization scheme governed by the linear system
${\bf L}{\bf y}={\bf x},$ (1)
where $\mathbf{y}\in\mathbb{R}^{K}$ is the current pattern to be optimized, $\mathbf{x}\in\mathbb{R}^{N}$ is the discretized volume current density vector, and $\mathbf{L}\in\mathbb{R}^{N\times K}$ is the lead field matrix defining a linear forward mapping from current patterns to the resulting current density field. This system is derived by component-wise splitting of the full lead field matrix $\hat{\mathbf{L}}\in\mathbb{R}^{N\times K}$ and the full current field $\hat{\mathbf{x}}\in\mathbb{R}^{N}$, such that
$\hat{\mathbf{L}}=\begin{pmatrix}\hat{\mathbf{L}}_{1}\\ \hat{\mathbf{L}}_{2}\end{pmatrix},\quad\hat{\mathbf{x}}=\begin{pmatrix}\hat{\mathbf{x}}_{1}\\ \mathbf{0}\end{pmatrix},$ (2)
where $\hat{\mathbf{x}}_{1}\in\mathbb{R}^{N_{1}}$ corresponds to the target region of interest and the zero component enforces field suppression elsewhere. The reduced system (1) is then defined by
$\mathbf{L}=\begin{pmatrix}\mathbf{L}_{1}\\ \mathbf{L}_{2}\end{pmatrix}=\begin{pmatrix}\mathbf{P}\hat{\mathbf{L}}_{1}\\ \hat{\mathbf{L}}_{2}\end{pmatrix},\quad\mathbf{x}=\begin{pmatrix}\mathbf{x}_{1}\\ \mathbf{0}\end{pmatrix}=\begin{pmatrix}\mathbf{P}\hat{\mathbf{x}}_{1}\\ \mathbf{0}\end{pmatrix}.$ (3)
Here, $\mathbf{P}\in\mathbb{R}^{N_{1}\times N_{1}}$ is a projection matrix that projects vectors onto the direction of $\hat{\mathbf{x}}_{1}$, which is assumed to have fixed direction. This projection ensures that alignment between the induced and target fields is measured only along the direction of interest. In this formulation, the upper part $\hat{\mathbf{L}}_{1}\mathbf{y}=\hat{\mathbf{x}}_{1}$ defines the focused current field in the target region, while the lower part $\hat{\mathbf{L}}_{2}\mathbf{y}=\mathbf{0}$ defines the nuisance field to be suppressed outside the region of interest.
A.1 L1-norm regularized L1-norm fitting
The general form of the L1-norm regularized L1-norm fitting (L1L1) method described by Galaz Prieto et. al [32, 57] is an optimization problem designed to minimize a linear objective function of continuous real variables subject to linear constraints. The optimization problem needs to find the best matching between ${\bf y}$, and the focused field via ${\bf L}{\bf y}={\bf x}$. The task is
[…]
@@ A.1 L1-norm regularized L1-norm fitting
The linear constraint 5 implies that the individual current injection for each $\ell$-th electrode is limited to a $\Gamma$ value; constraint 6 implies the total current injection flowing through the system is within a safety $\mu$-value limit; and constraint 7 indicates the total sum of electrical activity from every active electrode in the system must be equal to zero. The regularization parameter ${\bf\alpha}$ sets the L1-regularization level regarding a scaling value $\zeta=\|{\bf L}\|_{1}$. The function
@@
A.2 Tikhonov-Regularized Least Squares
In TLS estimation [37, 38] is formulated as the following optimization problem:

### v2

@@ 4 Discussion
This study explores the application of the recently introduced metaheuristic L1L1 method [galazprieto_2022_L1vsL2] to optimize current patterns and electrode configurations for DBS leads. The L1L1 method was applied in two variants, L1L1(A) constrains $\varepsilon$ to the interval $160,0]\,\text{dB}$ and L1L1(B) to $10,0]\,\text{dB}$ reflecting different levels of a priori modeling uncertainty [schmidt2016uncertainty, athawale2019statistical]. For comparison, optimization performance using Reciprocity Principle (RP)  [FERNANDEZCORAZZA2020116403] and Tikhonov Regularized Least Squares (TLS) [dmochowski2011optimized] methods were included. The core advantage in applying the L1L1 method is the capacity to incorporate a priori uncertainty through an explicit constraint, offering an advantage over traditional approaches, such as the least-square methods, which often do not account for such variability.
Within the optimization framework, the focused current density $\Gamma$, nuisance current density $\Xi$, and field ratio $\Theta=\Gamma/\Xi$ are implemented as the three metrics to measure spatial current distributions influenced by the quality of the forward modelling, position and orientation of the current dipole, and uncertainness which is defined through a threshold on the nuisance current density, determined by the relative attenuation of the lead field matrix with respect to a target stimulation site. This attenuation is quantified in decibels and corresponds to typical falloff patterns observed in the volume of tissue activated (VTA) [zhang2020steering].
In noiseless conditions, solutions using a more permissive threshold value (L1L1(A) case) often exhibited unrealistically high field ratio values, approximately fifty times greater than those derived using a more conservative threshold (L1L1(B) case). The addition of noise promoted the optimizer’s reliance from convergence towards over-idealized solutions, forcing the algorithm to select current patterns that remain effective even under slight model variations, producing more physiologically realistic distributions.
@@
The estimates for the set $\mathcal{V}_{\delta}$, where the lead field amplitude attenuates by a factor $\delta$ for $\delta\in[-40,-10]\,\text{dB}$, align closely with established knowledge regarding VTA. For $\delta\approx-40\,\text{dB}$, the model predicts attenuation within a spatial region consistent with a VTA radius of approximately 4.0 mm, in line with Butson and McIntyre [butson_2007_patient], Miocinovic et al. [miocinovic_2006_subthalamic], and Mädler and Coenen [madler2012activated]. This suggests that the noise levels used (PSNR 40 dB and 50 dB) represent reasonable bounds for physiological uncertainty. At PSNR 40 dB, the dynamic range of $\Theta$ is determined by the maximum lead field attenuation within the VTA, whereas PSNR 50 dB supports the assumption that each position within the VTA can be independently and coherently targeted.
@@
When comparing electrode configurations, increasing the number of contact points generally improves adaptability and control over the resulting electric field. Parallel configurations, i.e., alignment of the current dipole with the DBS lead, produced localized activation near the tip of the lead, whereas perpendicular configurations yielded a broader medially shifted targeting. This is driven by reduced distances between electrodes and target locations, better alignment of the induced field with the target current direction, and greater flexibility due to more DOFs.
@@
Overall, the findings provide a proof-of-concept for integrating a priori uncertainty modeling into numerically simulated DBS optimization. When sources of uncertainty, such as anatomical variability, conductivity inhomogeneities, or lead placement deviations, are known or estimated, the L1L1 method can improve focality and directionality of stimulation, enabling more selective targeting of desired structures while minimizing off-target activation. Such improvements have the potential to enhance therapeutic precision and reduce side effects in clinical practice [fasano2021experience].
@@ 4.1 Limitations and Future Work
While the current uncertainty modeling strategy provides a principled framework for constraining the optimization based on lead field attenuation, it also has several limitations. First, the threshold values used to define acceptable nuisance current levels are heuristic and may not fully capture the complexity of individual patient anatomies or tissue conductivities. This simplification assumes spatial uniformity in attenuation, which may not hold, for example, in the presence of anisotropic or heterogeneous tissue properties [dannhauer2011modeling]. Additionally, the approach relies on a static surrogate model of the VTA and does not directly incorporate a probabilistic model of uncertainty [schmidt2016uncertainty, athawale2019statistical], which would allow for a more comprehensive treatment of variability. Potential alternative surrogate approaches exist, such as, threshold-based curve fitting [butson_2006_electrodesign], and statistical emulators such as Gaussian process classifiers [DeLaPava_2017_accelerating, Orozco_2019_KernelDBS], which approach VTA in different ways.
The Gaussian white-noise model is adopted because it furnishes the maximally uninformative representation of uncertainty under a finite-variance constraint [Jaynes2003, CoverThomas2006]. Given that the present optimization framework does not specify a detailed probabilistic law for perturbations, this choice avoids introducing unsupported structure and provides a neutral, information-theoretic baseline. More sophisticated uncertainty quantification techniques, such as Markov Chain Monte Carlo (MCMC)-based Bayesian sampling [kaipio2006statistical] or deep-learning approaches to predictive uncertainty [schmidt2016uncertainty, athawale2019statistical], could be employed when richer data are available.
Compared to the reference methods, the L1L1 method is also slightly more demanding in terms of computing time. However, the increase is not significant and does not limit its practical use. Other mathematical optimization techniques, such as TLS, may offer promising alternatives alongside the L1L1 method for further refining the a priori uncertainty modeling adopted in this study. A logical direction for future research is the development of a posteriori models, where the current approach could be validated in more advanced experimental or clinical settings. This would enable a deeper understanding of the potential hyperparameter realizations in real-world applications. Translating these computational findings into clinical practice will require rigorous validation through empirical testing and clinical trials.
This manuscript does not address clinical implications; however, for the configuration studied, certain anatomical structures, such as the lamina medullaris near the modeled lead trajectory, may require special consideration in clinical practice to avoid stimulation-induced side effects. Translating these methodological results to clinical applications will necessitate examining the relationship between the L1L1 method and uncertainties in individual factors such as lead placement, tissue conductivity and anisotropy, and patient-specific anatomical variations including thalamic segmentation. To support this, we employ finite element (FE) discretizations incorporating thalamic subdivisions through a boundary-fitted multi-compartment FE mesh together with a Complete Electrode Model (CEM) [pursiainen2017forward], as in the present study.
@@ 5 Conclusions
This study represents an advancement in optimizing the volumetric current fields for DBS by leveraging the metaheuristic L1L1 method. Previous contact selection studies, such as [Anderson_2018] and later [Janson_2020_Activation], have used convex optimization to determine optimal stimulation settings and this represents at addressing varying levels of a priori lead field uncertainty, which can be attributed as patient anatomy, misleading lead placement, and tissue properties, which often vary in clinical practice.
Our findings demonstrate that the method can achieve relatively high field ratios with a carefully chosen activation threshold $\varepsilon$ value. This, in turn, has the potential to improve individualized stimulation outcomes. An ongoing work is to integrate patient-specific data, such as thalamic segmentation, to improve stimulation accuracy. This involves using boundary-fitted multi-compartment FE meshes [galazprieto_2023_mesh] and a Complete Electrode Model [pursiainen2017forward], as in this study.
@@ 7 Author Contributions
FGP: Conceptualization, Data curation, Formal analysis, Investigation, Methodology, Software, Visualization, Writing—original draft. AL: Conceptualization, Data curation, Investigation, Methodology, Project administration, Writing—original draft. MS: Formal analysis, Supervision, Validation, Writing—review & editing. SP: Conceptualization, Formal analysis, Funding acquisition, Investigation, Methodology, Project administration, Resources, Software, Supervision, Validation, Visualization, Writing—review & editing.
@@ 8 Funding
FGP, MS, and SP were supported by the Research Council of Finland (RCF) through the Center of Excellence in Inverse Modelling and Imaging 2018–2025 (RCF 353089), Exploratory Study for Radar Tomography of Dimorphos (RCF 359198), and the Flagship of Advanced Mathematics for Sensing, Imaging and Modelling (FAME) (RCF 359185); DAAD (German Academic Exchange Service) project ”Non-invasively reconstructing and inhibiting activity in focal epilepsy” (354976, 367453); and the ERA Personalised diagnosis and treatment for refractory focal paediatric and adult epilepsy (PerEpi) project (RCF 344712). MS holds Academy Research Fellowship 2025–2029 (RCF 371055).
9 Ethical Statement
This work is based exclusively on mathematical modeling and numerical simulations, using an openly available MRI dataset [piastra_maria_carla_2020_3888381] as input. No new data was collected, and no sensitive or personally identifiable information was processed at any stage of the study.
@@ Appendix A Mathematical Optimization Scheme
A.1 Formulation of the Lead Field Matrix
The linear forward mapping between electrode currents and the resulting volumetric current density distribution is given by the real-valued $(N+M)\times K$ lead field matrix $\mathbf{L}$:
$\mathbf{L}\,\mathbf{y}=\mathbf{x},$ (1)
where $\mathbf{y}\in\mathbb{R}^{K}$ denotes the applied current pattern across $K$ electrode contacts, and $\mathbf{x}\in\mathbb{R}^{N+M}$ is the discretized volumetric current density with each entry corresponding to a local source element parameterized by three Cartesian dipolar degrees of freedom (DOFs), representing current density along the $x$-, $y$-, and $z$-axes. For our optimization framework, $\mathbf{L}$ is partitioned into target and nuisance components:
$\mathbf{L}=\begin{pmatrix}\mathbf{L}_{1}\\ \mathbf{L}_{2}\end{pmatrix},\qquad\mathbf{x}=\begin{pmatrix}\mathbf{x}_{1}\\ \mathbf{0}\end{pmatrix}$, (2)
where $\mathbf{L}_{1}\in\mathbb{R}^{N\times K}$ models the target region, $\mathbf{L}_{2}\in\mathbb{R}^{M\times K}$ models the nuisance region, and $\mathbf{x}_{1}\in\mathbb{R}^{N}$ defines the desired field in the target region. To ensure that only the direction of $\mathbf{x}_{1}$ is prescribed, we introduce the projection
$\mathbf{P}=\frac{\mathbf{x}_{1}\mathbf{x}_{1}^{\mathsf{T}}}{\|\mathbf{x}_{1}\|_{2}^{2}}\in\mathbb{R}^{N\times N},$ (3)
which enforces alignment of the induced field with $\mathbf{x}_{1}$. The resulting reduced formulation is:
$\begin{pmatrix}\mathbf{P}\mathbf{L}_{1}\\ \mathbf{L}_{2}\end{pmatrix}\mathbf{y}=\begin{pmatrix}\mathbf{P}\mathbf{x}_{1}\\ \mathbf{0}\end{pmatrix}$. (4)
The first block enforces field orientation within the target region, while the second enforces suppression outside it.
A.2 L1-norm Regularized L1-norm Fitting (L1L1)
The general form of the L1-norm regularized L1-norm fitting (L1L1) method described by Galaz Prieto et. al [galazprieto_2022_L1vsL2] is a linear programming optimization problem designed to minimize a linear objective function of continuous real variables subject to linear constraints. The optimization problem needs to find the best matching between $\bf y$}, and the focused field via ${\bf L}{\bf y}={\bf x}$. The task is
$\underset{y}{\textrm{minimize}}$ $\quad\left\|\begin{pmatrix}{\bf L}_{1}{\bf y}-{\bf x}_{1}\\ \Psi_{\varepsilon}[\nu^{-1}{\bf L}_{2}{\bf y}]\end{pmatrix}\right\|_{1}+{\alpha}\zeta\|{\bf y}\|_{1},$ (5)
@@ A.2 L1-norm Regularized L1-norm Fitting (L1L1)
The linear constraint 6 implies that the individual current injection for each $\ell$-th electrode is limited to a $\gamma$ value; constraint 7 implies the total current injection flowing through the system is within a safety $\mu$-value limit; and constraint 8 indicates the total sum of electrical activity from every active electrode in the system must be equal to zero. The regularization parameter $\bf\alpha$} sets the L1-regularization level regarding a scaling value $\zeta=\|{\bf L}\|_{1}$. The function
@@
To evaluate stimulation performance and to quantify the trade-off between target and undesired current spread, we define three decision variables: the focused current density $\Gamma$, the nuisance current density $\Xi$, and the resulting field ratio $\Theta$.
The focused current density $\Gamma$ (A/m2) quantifies the degree of field alignment between the induced current distribution and the desired stimulation target. It is defined as
$\Gamma=\frac{{\bf x}_{1}^{T}{\bf L}_{1}{\bf y}}{\|{\bf x}_{1}\|_{2}},$ (10)
where ${\bf x}_{1}\in\mathbb{R}^{N}$ represents the target volumetric current density template specifying the spatial and directional characteristics of the desired field within the stimulation region. The inner product ${\bf x}_{1}^{T}{\bf L}_{1}{\bf y}$ measures how strongly the induced field ${\bf L}_{1}{\bf y}$ aligns with the target pattern, while normalization by $\|{\bf x}_{1}\|_{2}$ scales the result to yield a field intensity expressed in A/m2.
The nuisance current density $\Xi$ (A/m2) characterizes the average magnitude of the induced field in off-target regions and is defined by
$\Xi=\frac{\|{\bf L}_{2}{\bf y}\|_{2}}{\sqrt{M}}.$ (11)
Dividing the Euclidean norm $\|{\bf L}_{2}{\bf y}\|_{2}$ by $\sqrt{M}$ yields the root-mean-square (RMS) field amplitude per node,
$\frac{\|{\bf L}_{2}{\bf y}\|_{2}}{\sqrt{M}}=\sqrt{\frac{1}{M}\sum_{m=1}^{M}\big(({\bf L}_{2}{\bf y})_{m}\big)^{2}}.$
The division by $\sqrt{M}$ converts the Euclidean norm into the root-mean-square (RMS) field amplitude per node, thereby making $\Xi$ invariant to the total number of discretization points in the nuisance domain. In cases where only a subset of nodes exhibits appreciable current density, the normalization may alternatively be restricted to the active constraint support, i.e., replacing $M$ with the number of nodes satisfying $|({\bf L}_{2}{\bf y})_{m}|\geq\varepsilon\nu$.
The field ratio $\Theta$ is introduced to characterize the overall selectivity of stimulation:
$\Theta=\frac{\Gamma}{\Xi}.$ (12)
A high $\Theta$ value reflects both strong current focusing within the target and minimal activation of non-target regions.
A.3 Tikhonov-Regularized Least Squares
TLS estimation [dmochowski2011optimized] is formulated as the following optimization problem: